Find a basis for the null space, row space and column space of matrix
1
Expert's answer
2020-12-02T19:05:07-0500

Let the matrix is "A=\begin{bmatrix}
1 & 0 & 0 & 1\\
0 & 1 & 0 & 1 \\
0 & 0 & 1 & 1
\end{bmatrix}"


Basis for row space of "A=\begin{Bmatrix}
[1 & 0 & 0 & 1], [0 & 1 & 0 & 1],[0 & 0 & 1 & 1]
\end{Bmatrix}"


In the 1, 2 and 3, it contains pivot hence,

"=\{\begin{bmatrix}
1 & 0 & 0 \\
0 & 1 & 0 \\
0 & 0 & 1 
\end{bmatrix}\}"

If "Ax=0"

"\begin{bmatrix}
x_1\\
x_2 \\
x_3\\
x_4
\end{bmatrix}" "=\begin{bmatrix}
-1\\
-1 \\
-1\\
1
\end{bmatrix}"


Hence, it is much better for the null space "=\begin{bmatrix}
-1\\
-1 \\
-1\\
1
\end{bmatrix}"